# Latent Gaussian dynamic factor modeling and forecasting for multivariate count time series

### v2

@@ 6.1.2 Selection of the number of factors and lag order
Figure 2 depicts the frequencies of estimated $r$ in 100 replications. From the figure, PC method outperforms all baselines (IC1–3, ED) in selecting true $r$, and performs similarly to Fac. Interestingly, the traditional information criterion-based or scree plot-based approaches fail for this model. The quality of estimation by the BCV-based approaches follows the pattern in the estimation of model parameters. That is, as the dimension $d$ and sample length $T$ increase, so do the percentages of correctly estimated $r$. Another observation is that the larger number of factor series deteriorates the performance of cross-validations. This can be seen for both cross-validation schemes, which is also an expected phenomenon with the standard factor models. In terms of marginal distributions, cross-validation schemes work best for negative binomial, followed by Poisson, multinomial, and Bernoulli marginal distributions. This indicates that cross-validation schemes improve when marginal distributions tend to take a larger number of values so that they are more akin to continuous distributions.
Next, for the simulation of selecting the lag order, we fix $r=2$ but consider $p=1$ and $p=3$. The same setting as above is used for $p=1$, but the transition matrix $\Psi$ and covariance matrix $\Sigma_{\eta}$ are replaced by $\mbox{diag}(\Psi_{h})_{ii}=\rho_{h}$ with $\rho_{h}=0.2,0.3,0.4$ for $h=1,2,3$, and $\Sigma_{\eta}=0.3448I_{2}$, respectively. The four IC methods in (3.21) are considered as ad hoc baselines. Our approach is denoted by BCV. In addition, we also apply the cross-validation scheme described in Section 3.3 to both latent Gaussian series $\{Z_{t}\}$ and $\{Y_{t}\}$, by assuming that those series are observable. We denote them Low dim and Gaussian, respectively.
Figure 3 shows the frequencies of estimated $p$ in 100 replications. The IC methods perform poorly under all conditions. Although the BCV-based approaches also underperform, applying the cross-validation scheme to discrete-valued observed series performs quite similarly to applying it directly to both latent Gaussian series $\{Z_{t}\}$ and $\{Y_{t}\}$. Similar to the rank selection in Figure 2, as the dimension $d$ and the sample length $T$ increase, the performance appears to improve. However, this improvement seems relatively minor compared to the rank selection. Taking a much larger $T$ (e.g., $T=1,000$) was checked to yield predominantly the true lag order (with the plots omitted). Interestingly, unlike the different distributions of the estimated ranks depending on the marginal distributions, the estimated lag orders appear less sensitive to the marginal distributions of the observed series. Overall, a more accurate lag order selection procedure is still desirable, but note that this is even for the case when observing the $r-$dimensional series $\{Y_{t}\}$.
@@ 6.2 Forecasting
Table 3 reports the results from the observations $X_{t}$, in terms of the mean of RMSEs of forecasting errors and accuracy of forecasting including three benchmarks. Similarly to the two latent processes, the RMSEs of the forecasting errors increase as the forecasting horizon increases. Furthermore, the RMSEs of the forecasting errors for the observed series tend to be larger than those for the latent processes. It is interesting to note the larger forecasting errors for marginal distributions that take more values, such as negative binomial, followed by Poisson, multinomial, and Bernoulli distributions. This was not the case for the RMSEs of the forecasting errors of latent processes. Similar tendencies as for the RMSEs of the forecasting errors are observed with the accuracy measure. Finally, compared to Last, Marginal, and Null approaches, consistently better forecasting performance is noted for the proposed method (the larger accuracy is associated with better predictors).
@@ 7 Application
To demonstrate the utility of the proposed model, we consider individual-level time series consisting of daily self-report measures of personality collected by Borkenau and Ostendorf, (1998). That study was designed to explore items describing personal emotions representing the “Big Five” factors of personality. The structure of the data is as follows. 30 items of emotions where groups of 6 items are known to be related to one of the five factors in personality have been collected for 22 students over 90 days.
@@
All 30 items are thought to correspond to at least one of the “Big Five” factors. These categories are known to follow a factor structure, and are denoted as categories 1 through 5 below. Each evening the participants of the study were instructed to appraise their daily behavior for each item on a scale ranging from 0 to 6, with 6 indicating greater endorsement of the emotion that day. For illustrative purposes, we work with the data for one student out of the 22 available. The choice of the student was largely motivated by the following consideration: for many students, response on some items showed very little variability (i.e., being mostly constant) and such cases were not particularly interesting from the time series perspective. We reduce the effect of the two extreme observations 0 and 6 by merging them with 1 and 5, respectively, so that the new scale ranges from 1 to 5. The practitioners wishing to use our (or any other) model should first carry out some basic exploratory analysis of data, as we touch upon below.
@@
The data set and its time series plots are illustrated in Figure 4. Among 90 consecutive observations, we use the first 85 observations for estimation and the last 5 observations as a hold-out sample to evaluate our forecasts. The corresponding 6 items are grouped by the identified categories C1 through C5. The time series of individual items show substantial variability. We also see from the time plots that the dynamics of the 6 items in each category share common features. Hence, it is plausible to postulate the existence of a latent factor structure that drives the dynamics.
@@
The following remarks provide further evidence for the latent factor structure. Several estimated correlation matrices are depicted in Figure 5. The top panel presents the sample autocorrelation matrices of the observed series $X_{t}$. The bottom panel presents the estimated autocorrelation matrices of the latent series $Z_{t}$. For both panels, the left plot is for lag 0 and the right plot is for lag 1. One can see that both autocorrelation matrices for the same lag order are nearly indistinguishable. At lag 0, the plots in both panels show clear block patterns characteristic of the factor structure. Furthermore, the factor structure seems to be preserved through temporal dependence as suggested by the plots of the sample ACFs at lag 1, though it is less discernible.
@@
We work with the model assuming $r=5$ and $p=1$ for the sake of illustration. We also assume multinomial distributions on $\{1,2,3,4,5\}$ as marginal distributions. For the loadings matrices, we use the second identifiability condition (3.6). The left panel of Figure 6 presents the estimated loadings matrix and the right panel depicts the transition matrix of the factor series estimated through the method described in Section 3. This pattern is consistent with what we observe in Figure 5. Note that the transition matrix has relatively large values not only on the diagonal but also for some off-diagonal entries. The large off-diagonal values indicate that there are cross-correlations across the factor series.
@@
Figure 7 presents the generated particles $\{\widehat{Z}_{t}^{(k)}\}$ for 30 items within the last 5 observations and the predicted values of latent process $\{\widehat{Z}_{T+h|T}^{(k)}\}$ for next 5 forecasting steps, respectively. The items have been grouped using the 5 underlying categories. The three or four parallel lines in Figure 7 represent the thresholds, and each pair of lines represents a bin. To distinguish the values, we use different line types for each threshold. As explained above, the particles are generated at each time point to belong to a certain bin that matches the discrete observation for that dimension. This is why the particles stay within bins at all time points in the left panel. For some of the items, a thresholding line is placed outside the given vertical scale. This is due to the way the bins are defined. Note that since each bin is estimated through the observations, some values do not appear if they are not realized in the observation period. Those values are also excluded from the candidate forecasting values. On the other hand, since no further observations are assumed to be given after the observation period, the particles are generated by forecasting the latent process. For this reason, the particles in the right panel do not need to stay in the same bin. Note that all of the particles seem to converge to zero for the longer forecasting horizon. As explained in Section 5.3, this is natural when a stable VAR is used for forecasting. Note also that the particles are rather close to zero even for the first few horizons. This is a consequence of the interplay between the levels of signal (factors) and noise (errors) in the estimated model. The forecasted value naturally takes 0 when forecasting noise and thus downweights the factor forecast as the larger the noise magnitude is, since the variance of our latent process is 1 in each dimension.
Finally, Figure 8 shows the plots of the absolute differences between forecasts and the true values for each item. The items are ordered to match the categories expecting that each factor mainly affects the corresponding category. Overall, the proposed forecasting approach slightly outperforms the reference methods, by showing small absolute differences in general. In particular, as explained in Section 5.3, the forecasting performance becomes identical to the Marginal for the longer horizon. But for smaller horizons, our approach does better for 4 items, compared to 2 items doing better for Marginal. We naturally cannot draw overarching conclusions based on this evidence.
@@ 8 Conclusion
While our study advances a framework for latent Gaussian time series modeling of categorical observations collected over time, important questions remain. The question of how to employ time-varying covariates was discussed in Section 2.3 but left for future work. There are also potential improvements to make in terms of accuracy and computing time of our forecasting methods. Instead of employing standard particle filtering strategies, one could try other variants of sequential Monte Carlo sampling, for example, ensemble Kalman filtering in a high-dimensional regime (e.g., Katzfuss et al., , 2020).
Acknowledgement
MD was supported partially by the FAU Emerging Talents Initiative. VP was supported partially by the grants NSF DMS-1712966, DMS-2113662, and DMS-2134107.
@@ A.2 Auxiliary result and its proof
$\bm{\widehat{\Sigma}}_{X}$ $=\left(\frac{1}{N}\sum_{t=p+1}^{T}\widetilde{X}_{t-i}\widetilde{X}_{t-j}^{\prime}\right)_{i,j=1,\dots,p}$ (A.27)
$=\left(\frac{1}{N}\sum_{t=p+1}^{T}\left(G(Z_{t-i})-{\mathbb{E}}G(Z_{t-i})\right)\left(G(Z_{t-j})-{\mathbb{E}}G(Z_{t-j})\right)^{\prime}\right)_{i,j=1,\dots,p},$
@@
${\mathbb{E}}\bm{\widehat{\Sigma}}_{X}$ $=\left(\frac{1}{N}{\mathbb{E}}\sum_{t=p+1}^{T}\widetilde{X}_{t-i}\widetilde{X}_{t-j}^{\prime}\right)_{i,j=1,\dots,p}$ (A.28)
$=\left(\frac{1}{N}\sum_{t=p+1}^{T}({\mathbb{E}}(G(Z_{t-i})G(Z_{t-j})^{\prime})-{\mathbb{E}}G(Z_{t-i}){\mathbb{E}}G(Z_{t-j})^{\prime})\right)_{i,j=1,\dots,p}=\bm{\Sigma}_{X}.$
@@
The purpose of this section is to verify Assumption E.3. Recall that this assumption was critical in obtaining, for example, the relation (A.34). Rephrasing the assumption, the question is whether we can expect the spectral gap of the Markov chain (A.31) underlying our latent dynamic factor model to be bounded away from $1$ as the dimension $d$ is increasing. The following lemma formally states one possible quite general scenario under which Assumption E.3 is satisfied. Recall $\Psi_{1}=:\Psi$ and $\{\eta_{t}\}$ from Assumption 4.10.
@@ A.4.1 Preliminaries: Hermite polynomials
Let $\{X_{t}\}$ be a $D-$dimensional Gaussian process with zero mean and positive-definite covariance matrix $\Sigma_{X}$. We further introduce the eigenvectors $v_{1},\dots,v_{D}$ and eigenvalues $\lambda_{1},\dots,\lambda_{D}$ of $\Sigma_{X}$ such that $\Sigma_{X}v_{i}=\lambda_{i}v_{i}$. Define the linear functionals
@@ A.4.2 Preliminaries: Isonormal Gaussian processes
Let $\{X_{t}\}$ be a $D-$dimensional Gaussian process with zero mean and positive definite covariance matrix $\Sigma_{X}$. Let $v_{1},\dots,v_{D}$ denote the eigenvectors of $\Sigma_{X}$ and recall the linear functionals $W_{v_{i}}(\cdot)$ from (A.38). Then, the autocorrelations can be written as ${\mathbb{E}}W_{v_{i}}(X_{t})W_{v_{j}}(X_{s})=\rho_{ij}(t-s)=\mathop{\textrm{Corr}}(X_{i,t},X_{j,s})$. A Gaussian process $\{X_{t}\}$ can always be regarded as a subset of an isonormal Gaussian process $\{W(u):~{}u\in\mathfrak{H}\}$, where $\mathfrak{H}$ is a separable Hilbert space with scalar product $\langle\cdot,\cdot\rangle_{\mathfrak{H}}$. For every $m\geq 1$, we write $\mathfrak{H}^{\otimes m}$ to indicate the $m$th tensor power of $\mathfrak{H}$; $\mathfrak{H}^{\odot m}$ indicates the $m$th symmetric tensor power of $\mathfrak{H}$, equipped with the norm $\sqrt{m}\|\cdot\|_{\mathfrak{H}^{\otimes m}}$. We can assume that there exists $u_{i,t}\in\mathfrak{H}$ such that
@@
Recall the representation of our latent factor model in terms of a VAR(1) model from (A.31). From the representation (A.31), we know that $\{\mathcal{V}_{t}\}$ is a Markov chain. Then, setting $p=1$ and using (2.3), we can also write $\mathcal{V}_{t}$ as
@@
that itself is a $(r+d)$-dimensional Markov chain with ${\mathbb{E}}\zeta_{t}\zeta_{t}^{\prime}=:\operatorname{diag}(\Sigma_{\eta},\Sigma_{\varepsilon})$. For simplicity, we assume throughout the remainder of the proof that $\Sigma_{\varepsilon}=I_{d}$. Recall the definition of the right spectral gap from (4.13). Therefore, it suffices to bound the lag one correlation for functionals of the Markov chain (A.44). We use the following relationship:
@@
In order to bound (A.46), we study the autocovariance matrices at lags 00 and $1$ of $\{\mathcal{W}_{t}\}$ in (A.45). The autocovariance matrices of $\{\mathcal{W}_{t}\}$ for general lag $h$ are given by
@@
$\Gamma_{\mathcal{W}}(0)=\begin{pmatrix}\sum_{i=0}^{\infty}\Psi^{i}\Sigma_{\eta}\Psi^{i}&0_{r\times d}\\ 0_{d\times r}&I_{d}\end{pmatrix}=:\begin{pmatrix}\Sigma_{Y}&0_{r\times d}\\ 0_{d\times r}&I_{d}\end{pmatrix}=:\Sigma.$ (A.48)

### v3

@@ 6.1.2 Selection of the number of factors
Figure 2 depicts the frequencies of estimated $r$ in 100 replications. From the figure, PC method outperforms all baselines (IC1–3, ED) in selecting true $r$, and performs similarly to Fac. Interestingly, the traditional information criterion-based or scree plot-based approaches fail for this model. The quality of estimation by the BCV-based approaches follows the pattern in the estimation of model parameters. That is, as the dimension $d$ and sample length $T$ increase, so do the percentages of correctly estimated $r$. Another observation is that the larger number of factor series deteriorates the performance of cross-validations. This can be seen for both cross-validation schemes, which is also an expected phenomenon with the standard factor models. In terms of marginal distributions, cross-validation schemes work best for negative binomial, followed by Poisson, categorical, and Bernoulli marginal distributions. This indicates that cross-validation schemes improve when marginal distributions tend to take a larger number of values so that they are more akin to continuous distributions.
@@ 6.2 Forecasting
Table 3 reports the results from the observations $X_{t}$, in terms of the mean of RMSEs of forecasting errors and accuracy of forecasting including three benchmarks. Similarly to the two latent processes, the RMSEs of the forecasting errors increase as the forecasting horizon increases. Furthermore, the RMSEs of the forecasting errors for the observed series tend to be larger than those for the latent processes. It is interesting to note the larger forecasting errors for marginal distributions that take more values, such as negative binomial, followed by Poisson, categorical, and Bernoulli distributions. This was not the case for the RMSEs of the forecasting errors of latent processes. Similar tendencies as for the RMSEs of the forecasting errors are observed with the accuracy measure. Finally, compared to Last, Marginal, and Null approaches, consistently better forecasting performance is noted for the proposed method (the larger accuracy is associated with better predictors).
@@ 7 Application
To demonstrate the utility of the proposed model, we consider individual-level time series consisting of daily self-reported measures of personality, collected by Borkenau and Ostendorf, (1998). That study was designed to explore items describing personal emotions that represent the “Big Five” personality factors. The structure of the data is as follows. Time series for 30 emotion-related items were collected, where groups of 6 items are known to correspond to one of the five personality factors. These data were collected for 22 students over 90 days.
@@
All 30 items are believed to correspond to at least one of the “Big Five” personality factors. These factors are known to follow a structured categorization, denoted as categories 1 through 5 below. Each evening, participants in the study were instructed to evaluate their daily behavior for each item on a scale from 0 to 6, with 6 indicating the strongest endorsement of the emotion that day. For illustrative purposes, we focus on the data from one student out of the 22 available. This choice of the student was largely motivated by the following consideration: for many students, responses to some items exhibited very little variability (i.e., remained mostly constant), making them less interesting from a time series modeling perspective. To mitigate the influence of extreme observations 0 and 6, we merged them with 1 and 5, respectively, resulting in a new scale ranging from 1 to 5. Practitioners intending to use our model (or any other) should first conduct similarly basic exploratory data analysis.
@@
The data set and its time series plots are presented in Figure 3. Among 90 consecutive observations, we use the first 85 observations for estimation and reserve the last 5 as a holdout sample to evaluate our forecasts. The corresponding 30 items are grouped by the identified categories C1 through C5, each containing 6 items. The time series of individual items exhibit substantial variability. In addition, the time plots reveal that the dynamics of the 6 items within each category share common patterns. This suggests the plausible existence of a latent factor structure underlying these dynamics.
@@
The following remarks provide further evidence for the latent factor structure. Several estimated correlation matrices are depicted in Figure 4. The top panel presents the sample autocorrelation matrices of the observed series $\{X_{t}\}_{t=1,\ldots,85}$, while the bottom panel shows the estimated autocorrelation matrices of the latent series $\{Z_{t}\}_{t=1,\ldots,85}$. In both panels, the left plot corresponds to lag 0, and the right plot corresponds to lag 1. One can see that both autocorrelation matrices for the same lag order are nearly indistinguishable. At lag 0, the plots in both panels show clear block patterns characteristic of the factor structure. Furthermore, the factor structure appears to be preserved through temporal dependence, as suggested by the plots of the sample ACFs at lag 1, though it is less discernible.
@@
We work with the model assuming $r=5$ and $p=1$ for the sake of illustration. We also assume categorical distributions on $\{1,2,3,4,5\}$ as marginal distributions. For the loadings matrices, we use the second identifiability condition (3.6). The left panel of Figure 5 presents the estimated loadings matrix, while the right panel depicts the transition matrix of the factor series, estimated using the method described in Section 3. This pattern is consistent with what we observe in Figure 4. Note that the transition matrix has relatively large values not only on the diagonal but also for some off-diagonal entries. The large off-diagonal values indicate that there are cross-correlations across the factor series.
For the model diagnostics, Figure 6 shows the PIT histograms for 30 individual items, following the procedure in Section 5.4. With the exception of a few items, the heights of the PIT histograms across 10 bins tend to be uniformly equal, aligning with the 0.1 relative frequencies indicated by the black dashed lines.
@@
Figure 7 presents the generated particles $\{\widehat{Z}_{t}^{(k)}\}_{t=1,\ldots,T}$ for 30 items within the last 5 observations $t=81,\ldots,85$ and the predicted values of latent process $\{\widehat{Z}_{T+h|T}^{(k)}\}_{h=1,\ldots,5}$ for the next 5 forecasting steps, respectively. The items are grouped into five underlying categories. The three or four parallel lines in Figure 7 represent the thresholds, with each pair of lines forming a bin. To distinguish the values, we use different line types for each threshold. As explained above, the particles are generated at each time point to belong to a specific bin that corresponds to the discrete observation for that dimension. This explains why the particles stay within bins at all time points in the left panel. For some of the items, a thresholding line extends beyond the given vertical scale due to the way the bins are defined. Since each bin is estimated through the observations, some values do not appear if they are not realized in the observation period. These values are also excluded from the candidate forecasting values. On the other hand, since no further observations are assumed to be given after the observation period, the particles are generated by forecasting the latent process. As a result, the particles in the right panel do not need to stay in the same bin. Note that all of the particles seem to converge to zero for the longer forecasting horizon. As discussed in Section 5.3, this behavior is expected when a stable VAR is used for forecasting. In addition, the particles are rather close to zero even for the first few horizons. This is a consequence of the interplay between the levels of signal (factors) and noise (errors) in the estimated model. The forecasted value naturally approaches 0 when forecasting noise, which in turn downweights the factor forecast as the noise magnitude increases, given that the variance of our latent process is 1 in each dimension.
Finally, Figure 8 shows the plots of the absolute differences between forecasts and the true values for each item. The items are arranged according to their categories, under the expectation that each factor primarily influences its corresponding category. Overall, the proposed forecasting approach slightly outperforms the reference methods, generally exhibiting small absolute differences. As discussed in Section 5.3, the forecasting performance becomes identical to the Marginal for the longer horizon. But for shorter horizons, our approach performs better for 4 items, whereas the Marginal method performs better for 2 items. While this suggests a potential advantage, it is not sufficient to draw overarching conclusions based on this evidence.
@@ 8 Conclusion
While our study advances a framework for latent Gaussian time series modeling of categorical observations collected over time, important questions remain. The question of how to employ time-varying covariates is discussed in Supplemental Material but left for future work. There are also potential improvements to make in terms of accuracy and computing time of our forecasting methods. Instead of employing standard particle filtering strategies, one could try other variants of sequential Monte Carlo sampling, for example, ensemble Kalman filtering in a high-dimensional regime (e.g., Katzfuss et al., , 2020).
Acknowledgment
MD was supported partially by the FAU Emerging Talents Initiative. VP was supported partially by the grants NSF DMS-1712966, DMS-2113662, and DMS-2134107. The authors also thank two anonymous Reviewers for their many helpful comments.
@@ A.2 Auxiliary result and its proof
$\bm{\widehat{\Sigma}}_{X}$ $=\left(\frac{1}{T-p}\sum_{t=p+1}^{T}\widetilde{X}_{t-i}\widetilde{X}_{t-j}^{\prime}\right)_{i,j=1,\dots,p}$ (A.27)
$=\left(\frac{1}{T-p}\sum_{t=p+1}^{T}\left(G(Z_{t-i})-{\mathbb{E}}G(Z_{t-i})\right)\left(G(Z_{t-j})-{\mathbb{E}}G(Z_{t-j})\right)^{\prime}\right)_{i,j=1,\dots,p},$
@@
${\mathbb{E}}\bm{\widehat{\Sigma}}_{X}$ $=\left(\frac{1}{T-p}{\mathbb{E}}\sum_{t=p+1}^{T}\widetilde{X}_{t-i}\widetilde{X}_{t-j}^{\prime}\right)_{i,j=1,\dots,p}$ (A.28)
$=\left(\frac{1}{T-p}\sum_{t=p+1}^{T}({\mathbb{E}}(G(Z_{t-i})G(Z_{t-j})^{\prime})-{\mathbb{E}}G(Z_{t-i}){\mathbb{E}}G(Z_{t-j})^{\prime})\right)_{i,j=1,\dots,p}=\bm{\Sigma}_{X}.$
@@
The purpose of this section is to verify Assumption E.3. Recall that this assumption was critical in obtaining, for example, the relation (A.34). Rephrasing the assumption, the question is whether we can expect the spectral gap of the Markov chain (A.31) underlying our latent dynamic factor model to be bounded away from $1$ as the dimension $d$ is increasing. The following lemma formally states one possible quite general scenario under which Assumption E.3 is satisfied. Recall $\Psi_{1}=\Psi$ and $\{\eta_{t}\}$ from Assumption 4.10.
@@ A.4.1 Preliminaries: Hermite polynomials
Let $\{X_{t}\}_{t\in\mathbb{Z}}$ be a $D-$dimensional Gaussian process with zero mean and positive-definite covariance matrix $\Sigma_{X}$. We further introduce the eigenvectors $v_{1},\dots,v_{D}$ and eigenvalues $\lambda_{1},\dots,\lambda_{D}$ of $\Sigma_{X}$ such that $\Sigma_{X}v_{i}=\lambda_{i}v_{i}$. Define the linear functionals
@@ A.4.2 Preliminaries: Isonormal Gaussian processes
Let $\{X_{t}\}_{t\in\mathbb{Z}}$ be a $D-$dimensional Gaussian process with zero mean and positive definite covariance matrix $\Sigma_{X}$. Let $v_{1},\dots,v_{D}$ denote the eigenvectors of $\Sigma_{X}$ and recall the linear functionals $W_{v_{i}}(\cdot)$ from (A.38). Then, the autocorrelations can be written as ${\mathbb{E}}W_{v_{i}}(X_{t})W_{v_{j}}(X_{s})=\rho_{ij}(t-s)=\mathop{\textrm{Corr}}(X_{i,t},X_{j,s})$. A Gaussian process $\{X_{t}\}_{t\in\mathbb{Z}}$ can always be regarded as a subset of an isonormal Gaussian process $\{W(u):~{}u\in\mathfrak{H}\}$, where $\mathfrak{H}$ is a separable Hilbert space with scalar product $\langle\cdot,\cdot\rangle_{\mathfrak{H}}$. For every $m\geq 1$, we write $\mathfrak{H}^{\otimes m}$ to indicate the $m$th tensor power of $\mathfrak{H}$; $\mathfrak{H}^{\odot m}$ indicates the $m$th symmetric tensor power of $\mathfrak{H}$, equipped with the norm $\sqrt{m}\|\cdot\|_{\mathfrak{H}^{\otimes m}}$. We can assume that there exists $u_{i,t}\in\mathfrak{H}$ such that
@@
Recall the representation of our latent factor model in terms of a VAR(1) model from (A.31). From the representation (A.31), we know that $\{\mathcal{V}_{t}\}_{t\in\mathbb{Z}}$ is a Markov chain. Then, setting $p=1$ and using (2.3), we can also write $\mathcal{V}_{t}$ as
@@
that itself is a $(r+d)$-dimensional Markov chain with ${\mathbb{E}}\zeta_{t}\zeta_{t}^{\prime}:=\operatorname{diag}(\Sigma_{\eta},\Sigma_{\varepsilon})$. For simplicity, we assume throughout the remainder of the proof that $\Sigma_{\varepsilon}=I_{d}$. Recall the definition of the right spectral gap from (4.13). Therefore, it suffices to bound the lag one correlation for functionals of the Markov chain (A.44). We use the following relationship:
@@
In order to bound (A.46), we study the autocovariance matrices at lags 00 and $1$ of $\{\mathcal{W}_{t}\}$ in (A.45). The autocovariance matrices of $\{\mathcal{W}_{t}\}_{t\in\mathbb{Z}}$ for general lag $h$ are given by
@@
$\Gamma_{\mathcal{W}}(0)=\begin{pmatrix}\sum_{i=0}^{\infty}\Psi^{i}\Sigma_{\eta}\Psi^{i}&0_{r\times d}\\ 0_{d\times r}&I_{d}\end{pmatrix}=\begin{pmatrix}\Sigma_{Y}&0_{r\times d}\\ 0_{d\times r}&I_{d}\end{pmatrix}=\Sigma.$ (A.48)